# Onset of Planet Formation and the Transition Disk Phase

### v1

@@ Abstract
1 Introduction
@@
Our goal in this study is to examine how disk properties and their evolution, i.e., changes in disk structure caused by declining accretion rates, affect the trapping of pebbles at the DZIB, which is needed for the onset of IOPF. In §2 we introduce the physical model of our evolving protoplanetary disk and discuss the basic physics of pebble radial drift and trapping. We then examine the particular conditions for trapping at the DZIB, especially constraints on the size distribution of pebbles, and discuss the mass fraction of pebbles that is expected to be trapped at this location. In §3 we summarize our conclusions and discuss the implications of these results for the onset of planet formation in the IOPF scenario, as well as the emergence of the transition disk phase of protoplanetary disks.
2 Physical Model
2.1 Disk Structure
@@ 2.2 Pebble Drift
In the inner part of the protoplanetary disk, the midplane temperature is dominated by viscous heating powered by accretion, so the gas radial velocity is (Paper IV):
@@
where $\gamma\equiv 1.4\gamma_{1.4}$ with fiducial normalization to a value of 1.4 for $\rm H_{2}$ with rotational modes excited, $\kappa\equiv\kappa_{10}10\ {\rm cm^{2}\>g^{-1}}$ is the disk mean opacity (with fiducial normalization here appropriate for inner disk conditions; but, note we will more generally use tabulated opacities from Zhu et al., 2009), $\sigma_{\rm SB}$ is the Stefan-Boltzmann constant, $m_{*}\equiv m_{*,1}M_{\odot}$ is the stellar mass with fiducial normalization of $1\>M_{\odot}$, $f_{r}\equiv 1-\sqrt{r_{*}/r}$, $r_{*}$ is the stellar radius, and $r_{\rm au}\equiv r/(1\>{\rm au})$.
The disk aspect ratio, $h/r=c_{s}/v_{K}$, is an important metric of disk structure that can be related to pebble radial drift. It can be written in the form (Paper IV):
@@
2.3 Pebble Trapping at the DZIB
For the fastest drifting pebbles ($\tau_{\rm fric}=1$) at 0.1 au, when using all fiducial parameters, i.e., we have that $k_{P}=51/20=2.55$ in a disk of constant $\alpha$ and $\kappa$, then we find that $v_{\rm drift}\simeq 2.04\times 10^{4}k_{P}\alpha_{-4}^{-1/5}{\rm cm/s}$ and $v_{\rm drag}\simeq 0.97\alpha_{-4}^{4/5}{\rm cm/s}$. While these two velocity scales differ by several orders of magnitude, within the pressure maxima $k_{P}\simeq 0$, so that drift speeds become very small. As a result of inward gas motions, the pebble radial drift in fact needs to be outwards with respect to the gas, i.e., where $k_{P}$ is just becoming slightly smaller than zero. Furthermore, the value of $\alpha$ in this region is expected to be changing quite rapidly as one leaves the dead zone. The opacities may also be varying, e.g., due to the beginnings of dust destruction. The assumption of $\tau_{\rm fric}=1$ will also be invalid for the majority of pebbles, as gas densities are high and pebbles are more coupled with gas flow. With the disk profiles and drift equations, we can numerically calculate the minimum pebble size to be trapped at DZIB.
@@ 2.3 Pebble Trapping at the DZIB
Figure 3: (a) Left: Evolution of disk and planet properties as a function of accretion rate. Note, the evolution during planet formation is expected to be from a high accretion rate state (on the right) to a low accretion rate state (on the left), as indicated by the arrow at the bottom of the figure. These results assume constant opacity $\kappa=10\>{\rm cm^{2}\>g}^{-1}$ and $\alpha=10^{-4}$. Upper panel: Minimum size for a pebble to be trapped at the DZIB versus accretion rate (white solid line). Pebbles in the blue region are trapped, while those in the red region are dragged inward through the DZIB along with gas accretion. The black solid line is the critical size to enter the Stokes drag regime (above the line) from the Epstein drag regime (below the line). Middle panel: Location of pressure maximum predicted by Eq. 4. Lower panel: Mass of the first planet to form in the IOPF model due to shallow gap opening leading to pebble isolation (Eq. 2.5). (b) Right: As (a), but now shown as a function of radial location of the DZIB, which evolves from large to small distances, and with the middle panel showing accretion rate versus $r_{\rm DZIB}$.
From Equation 1, inside the pressure maximum of the DZIB, where gas orbits at super-Keplerian speeds, there is a competition between the tail wind induced outward radial drift and inward radial gas drag. For pebbles to be trapped, they need to be larger than a certain size to become decoupled from the inward radial gas drag and start to experience significant outward motion due to the super-Keplerian gas tail wind. The location of the DZIB, i.e., where the disk reaches a temperature of about 1,200 K, is estimated as (Papers I & IV):
@@
where $\rho_{p,3}$ is the pebble density normalized to $3\>{\rm g\>cm}^{-3}$. Note, here we have adopted the Epstein drag regime, which is shown to be valid (see below). Equation (5) includes a weak dependence on opacity. However, this can be approximated as a constant, since it takes values $\kappa=7.93,8.33,9.11\>{\rm cm^{2}\>g^{-1}}$ when $\dot{m}=10^{-8},10^{-9},10^{-10}M_{\odot}\>{\rm yr}^{-1}$, respectively. The main result is that the minimum pebble size required for trapping at the DZIB shows a dependence on disk accretion rate as $a_{p,{\rm trap}}\propto\dot{m}^{1/3}$. Thus, as the disk evolves and the accretion rate declines, a larger and larger fraction of pebbles is expected to be trapped at the DZIB.
@@
The top panels of Figure 3 show the dependence of $a_{p,{\rm trap}}$ as a function of accretion rate and as a function of radius of the DZIB. Note, the accretion rate is expected to decline as the disk evolves and, associated with this, $r_{\rm DZIB}$ is also expected to decrease. These panels also show the size of pebbles that would be subject to the Stokes drag regime, thus validating the use of the Epstein regime in the derivation of Eq. (5). Other panels in Figure 3 show how accretion rate and $r_{\rm DZIB}$ vary with each other and how the expected planet mass formed at the DZIB in the IOPF model (set by shallow gap opening) varies with these quantities.
2.4 Pebble trapping fraction at the DZIB
@@
Simulations with the DustPy code (Stammler & Birnstiel, 2022) of dust evolution within a fixed background gas model of the Paper IV disk, i.e., with $\dot{m}_{-9}=1$ and $\alpha=10^{-4}$, have been performed by Hjält (2024), who also find a narrow distribution of sizes of pebbles inside 1 au that is peaked at $\sim 0.6\>$cm (for fragmentation velocity of 1 m/s).
@@
In Figure 4a we show the evolution of the minimum size of pebbles to be trapped at the DZIB, $a_{\rm trap}$, as a function of disk accretion rate. We also indicate the peak pebble sizes of three assumed log-normal distributions, peaking at 0.5, 1, and 2 mm and with dispersion $\sigma_{p}=0.2$ dex, which are guided by the results of IOPF Paper IV. In Figure 4b, we then show the evolution of the resulting pebble trapping fraction at the DZIB, $f_{\rm trap}$. We see that this fraction rises from near zero during the early, high accretion rate stages, to approach unity by the final low accretion rate phase. The precise value of when this transition occurs, e.g., when $f_{\rm trap}=0.5$, depends on the assumed pebble size distribution. For the values of 0.5, 1 and 2 mm, these transition accretion rates are $\sim 6\times 10^{-10}$, $6\times 10^{-9}$, and $6\times 10^{-8}\>M_{\odot}\>{\rm yr}^{-1}$, respectively. We thus see that the onset of pebble trapping at the DZIB is quite sensitive to the pebble size distribution in this location, which thus motivates the need for improved studies of the processes of pebble growth and destruction in these environments.
Here, we discuss some additional considerations related to the pebble size distribution in DZIB regions. One process that may set a maximum size of pebbles is collisional fragmentation. The turbulent fragmentation limit on pebble radius is (Birnstiel et al., 2012):
@@
2.5 Impact of planet formation time
The planetary mass required for pebble isolation via shallow gap opening, i.e., the mass scale of planets forming via IOPF, is (Paper IV):
@@ 2.5 Impact of planet formation time
Thus, we see that the innermost planet’s mass varies from $M_{p,1}=1.18M_{\oplus}$ for $\dot{m}=10^{-10}\>M_{\odot}\>{\rm yr}^{-1}$, to $4.25M_{\oplus}$ for $\dot{m}=10^{-9}\>M_{\odot}\>{\rm yr}^{-1}$, to $15.3M_{\oplus}$ for $\dot{m}=10^{-8}\>M_{\odot}\>{\rm yr}^{-1}$ and that the required gap opening mass is significantly larger in the earlier, higher accretion rate phases when the DZIB is further out.
@@
$t_{\rm form,1}$ $=$ $\frac{M_{p,1}}{f_{\rm trap}\dot{m}_{p}}$ (12)
$=$ $1.28\times 10^{6}\frac{\gamma_{1.4}^{35/36}\kappa_{10}^{5/18}}{f_{\rm trap}f_{p}f_{s,0.01}}\alpha_{-4}^{2/9}m_{*,1}^{1/6}\frac{f_{r}^{5/9}}{\dot{m}_{-9}^{4/9}}\>{\rm yr}.$
Thus, we see that even when the trapping fraction rises to unity the timescale to form the first planet can be $\sim 1\>$Myr, which is a characteristic timescale of disk accretion rate evolution. However, as discussed in Paper IV, there could be a phase of boosted pebble supply rate during early phases of disk evolution, triggered by the time when a significant population of macroscopic pebbles develops that is able to engage in efficient sweep-up growth of the small grain dust population. Typical boost factors in the pebble supply rate in the first few $\times 10^{5}\>$yr were found to be $f_{p,{\rm boost}}\sim 10$. In the case of significant enhancements in pebble supply rates, the timescales implied by equation (12) should be regarded as upper limits.
@@
Finally, given the adopted disk evolutionary timescale, i.e., accretion rate declining by a factor of 10 every 1 Myr, and the particular values of $M_{p,1}$, there is some evolution in the accretion rate and thus $r_{\rm DZIB}$ and $M_{p,1}$ during the time when most of the planetary mass is being gathered. Figure 5 indicates (solid circles for $f_{p,{\rm boost}}=1$; open circles for $f_{p,{\rm boost}}=10$) the locations along the evolutionary tracks when the time integrated trapped pebble mass equals the required planetary gap opening mass at the current DZIB location.

### v2

@@ Abstract
show]jctan.astro@gmail.com
I Introduction
@@
Our goal in this study is to examine how disk properties and their evolution, i.e., changes in disk structure caused by declining accretion rates, affect the trapping of pebbles at the DZIB, which is needed for the onset of IOPF. In §II we introduce the physical model of our evolving protoplanetary disk and discuss the basic physics of pebble radial drift and trapping. We then examine the particular conditions for trapping at the DZIB, especially constraints on the size distribution of pebbles, and discuss the mass fraction of pebbles that is expected to be trapped at this location. In §III we summarize our conclusions and discuss the implications of these results for the onset of planet formation in the IOPF scenario, as well as the emergence of the transition disk phase of protoplanetary disks.
II Physical Model
II.1 Disk Structure
@@ II.2 Pebble Drift
In the inner part of the protoplanetary disk, the midplane temperature is dominated by viscous heating powered by accretion, so the gas radial velocity is (Paper I, eq. 12, note the differences in normalization of $\alpha$ and $r_{\rm au}$):
@@
where $\gamma\equiv 1.4\gamma_{1.4}$ with fiducial normalization to a value of 1.4 for $\rm H_{2}$ with rotational modes excited, $\mu=2.33m_{\rm H}$ is the mean molecular mass, $\kappa\equiv\kappa_{10}10\ {\rm cm^{2}\>g^{-1}}$ is the disk mean opacity (with fiducial normalization here appropriate for inner disk conditions; but, note we will more generally use tabulated opacities from Zhu et al., 2009), $\sigma_{\rm SB}$ is the Stefan-Boltzmann constant, $m_{*}\equiv m_{*,1}M_{\odot}$ is the stellar mass with fiducial normalization of $1\>M_{\odot}$, $f_{r}\equiv 1-\sqrt{r_{*}/r}$, $r_{*}$ is the stellar radius, and $r_{\rm 0.1au}\equiv r/(0.1\>{\rm au})$. The gas radial velocity increases with the accretion rate $\dot{m}$ and decreases with radius $r$, with both scalings following a $2/5$ power-law dependence.
The disk aspect ratio, $h/r=c_{s}/v_{K}$, is an important metric of disk structure that can be related to pebble radial drift. It can be written in the form (Paper III, eq. 30):
@@
The inner disk, where viscous heating sets the temperature structure, is quite thin, exhibiting a small flaring index of $0.05$.
@@
II.3 Pebble Trapping at the DZIB
For the fastest-drifting pebbles ($\tau_{\rm fric}=1$) at 0.1 au, adopting the fiducial parameters such that $k_{P}=51/20=2.55$ in a disk with constant $\alpha$ and $\kappa$, we obtain $v_{\rm drift}\simeq 2.04\times 10^{4}k_{P}\alpha_{-4}^{-1/5}{\rm cm\,s^{-1}}$ and $v_{\rm drag}\simeq 0.97\alpha_{-4}^{4/5}{\rm cm\,s^{-1}}$. Although these two characteristic velocity scales differ by several orders of magnitude, the drift velocity becomes negligible near pressure maxima where $k_{P}\simeq 0$. As a consequence of the inward motion of the gas, the radial drift of pebbles must be directed outward relative to the gas flow, occurring in regions where $k_{P}$ is slightly negative. Furthermore, the value of $\alpha$ in this region is expected to be changing quite rapidly as one leaves the dead zone. The opacities may also be varying, e.g., due to the beginnings of dust destruction. The assumption of $\tau_{\rm fric}=1$ will also be invalid for the majority of pebbles, as gas densities are high and pebbles are more coupled with gas flow. With the disk profiles and drift equations, we can numerically calculate the minimum pebble size to be trapped at DZIB.
@@ II.3 Pebble Trapping at the DZIB
Figure 3: (a) Left: Evolution of disk and planet properties as a function of accretion rate. Note, the evolution during planet formation is expected to be from a high accretion rate state (on the right) to a low accretion rate state (on the left), as indicated by the arrow at the bottom of the figure. These results assume constant opacity $\kappa=10\>{\rm cm^{2}\>g}^{-1}$ and $\alpha=10^{-4}$. Upper panel: Minimum size for a pebble to be trapped at the DZIB versus accretion rate (white solid line). Pebbles in the blue region are trapped, while those in the red region are dragged inward through the DZIB along with gas accretion. The black solid line is the critical size to enter the Stokes drag regime (above the line) from the Epstein drag regime (below the line). Middle panel: Location of pressure maximum predicted by Eq. 4. Lower panel: Mass of the first planet to form in the IOPF model due to shallow gap opening leading to pebble isolation (Eq. II.5). (b) Right: As (a), but now shown as a function of radial location of the DZIB, which evolves from large to small distances, and with the middle panel showing accretion rate versus $r_{\rm DZIB}$.
From Equation 1, inside the pressure maximum of the DZIB, where gas orbits at super-Keplerian speeds, there is a competition between the tail wind induced outward radial drift and inward radial gas drag. For pebbles to be trapped, they need to be larger than a certain size to become decoupled from the inward radial gas drag and start to experience significant outward motion due to the super-Keplerian gas tail wind. The location of the DZIB, i.e., where the disk reaches a temperature of about 1,200 K, is estimated as (Paper I, eq. 11 & Paper IV, eq. 13):
@@
where $\rho_{p,3}$ is the pebble density normalized to $3\>{\rm g\>cm}^{-3}$. Note, here we have adopted the Epstein drag regime. The midplane density at 0.1 au is (Paper I, eq. 2 and Paper III, eq. 31):
$\rho_{g}$ $=$ $2.52\times 10^{-8}\gamma_{1.4}^{-6/5}\kappa_{10}^{-3/10}\alpha_{-4}^{-7/10}$ (6)
$\times$ $m_{*,1}^{11/20}(f_{r}\dot{m}_{-9})^{2/5}r_{\rm 0.1au}^{-33/20}\>\rm g\>cm^{-3}$
The mean free path $\lambda=1/({n_{\rm H2}}{\sigma_{\rm H2}})$, with the number density $n_{\rm H2}=\rho_{g}/\mu$ and the cross section of the molecule $\rm H_{2}$} $\sigma_{\rm H2}\>=\>2\times 10^{-15}\rm cm^{2}$, so we have the upper limit on the size of pebbles in the Epstein regime (Paper I, eq. 10):
$\frac{9}{4}\lambda$ $=$ $0.173\gamma_{1.4}^{6/5}\kappa_{10}^{3/10}\alpha_{-4}^{7/10}$ (7)
$\times$ $m_{*,1}^{-11/20}(f_{r}\dot{m}_{-9})^{-2/5}r_{\rm 0.1au}^{33/20}\>\rm cm$
Note that this upper limit scales with $r^{33/20}$, so at $r_{\rm DZIB}$, it increases to $\sim 0.6\>$cm, an order of magnitude greater than $a_{p,{\rm trap}}$. More complete comparisons between these two sizes are shown in the top panels of Figure 3.
Equation (5) includes a weak dependence on opacity. However, this can be approximated as a constant, since it takes values $\kappa=7.93,8.33,9.11\>{\rm cm^{2}\>g^{-1}}$ when $\dot{m}=10^{-8},10^{-9},10^{-10}M_{\odot}\>{\rm yr}^{-1}$, respectively. The main result is that the minimum pebble size required for trapping at the DZIB shows a dependence on disk accretion rate as $a_{p,{\rm trap}}\propto\dot{m}^{1/3}$. Thus, as the disk evolves and the accretion rate declines, a larger and larger fraction of pebbles is expected to be trapped at the DZIB.
@@
The top panels of Figure 3 show the dependence of $a_{p,{\rm trap}}$ as a function of accretion rate and as a function of radius of the DZIB. Note, the accretion rate is expected to decline as the disk evolves and, associated with this, $r_{\rm DZIB}$ is also expected to decrease. These panels also show the size of pebbles that would be subject to the Stokes drag regime, thus validating the use of the Epstein regime in the derivation of Eq. (5). Other panels in Figure 3 show how accretion rate and $r_{\rm DZIB}$ vary with each other and how the expected planet mass formed at the DZIB in the IOPF model (set by shallow gap opening) varies with these quantities. We note that the precise definition of the planet mass is introduced later in Eq. II.5.
II.4 Pebble trapping fraction at the DZIB
@@
Simulations with the DustPy code (Stammler & Birnstiel, 2022) of dust evolution within a fixed background gas model of the Paper IV disk, i.e., with $\dot{m}_{-9}=1$ and $\alpha=10^{-4}$, have been performed by Hjält (2024), who also find that inside 1 au most of solids mass are dominated by the upper end of the pebble size, with a narrow distribution that is peaked at $\sim 0.6\>$cm (for fragmentation velocity of 1 m/s).
@@
In Figure 4a we show the evolution of the minimum size of pebbles to be trapped at the DZIB, $a_{\rm trap}$, as a function of disk accretion rate. We also indicate the peak pebble sizes of three assumed log-normal distributions, peaking at 0.5, 1, and 2 mm and with dispersion $\sigma_{p}=0.2$ dex, which are guided by the results of IOPF Paper IV. Note that this log-normal distribution is intended to approximate the upper end of the evolved pebble size distribution; the full width of the large-grain component that dominates the solid mass is approximately 1 dex (corresponding to $5\sigma$). In Figure 4b, we then show the evolution of the resulting pebble trapping fraction at the DZIB, $f_{\rm trap}$. We see that this fraction rises from near zero during the early, high accretion rate stages, to approach unity by the final low accretion rate phase. The precise accretion rate at which this transition occurs depends on the assumed pebble size distribution. For pebble sizes of 0.5, 1, and 2 mm, the transition, defined by $f_{\rm trap}=0.5$, occurs at accretion rates of $\sim 6\times 10^{-10}$, $6\times 10^{-9}$, and $6\times 10^{-8}\>M_{\odot}\>{\rm yr}^{-1}$, respectively. Keeping this log-normal size distribution, a smaller dispersion leads to a much sharper transition in trapping efficiency over the course of disk evolution. For example, adopting $\sigma=0.1$ dex with a peak size of 1 mm yields a minimum size of $\sim 0.5$ mm, in which case the trapping efficiency is nearly zero at $\dot{m}=10^{-8}\,M_{\odot}\,{\rm yr^{-1}}$ and then rises rapidly to almost 100% by $\dot{m}=10^{-9}\,M_{\odot}\,{\rm yr^{-1}}$. In contrast, a broader distribution with $\sigma=0.4$ dex corresponds to a “full width” spanning approximately 0.050.052020 mm. This wider range produces a trapping efficiency of $\sim 15\%$ even at $\dot{m}=10^{-6}\,M_{\odot}\,{\rm yr^{-1}}$, followed by a much flatter increase that does not approach unity until $\dot{m}\simeq 10^{-10}\,M_{\odot}\,{\rm yr^{-1}}$. We thus see that the onset of pebble trapping at the DZIB is quite sensitive to the pebble size distribution in this location, which thus motivates the need for improved studies of the processes of pebble growth and destruction in these environments.
Here, we discuss some additional considerations related to the pebble size distribution in DZIB regions. One process that may set a maximum size of pebbles is collisional fragmentation. The turbulent fragmentation limit on pebble radius is (Birnstiel et al., 2012, eq.8):
@@
II.5 Impact of planet formation time
The planetary mass required for pebble isolation via shallow gap opening, i.e., the mass scale of planets forming via IOPF, is (Paper IV, eq. 15):
@@ II.5 Impact of planet formation time
Note the differences from eq. 17 in Paper IV, where we replaced $r_{0.1,{\rm au}}$ with $r_{\rm DZIB}$. Because $M_{\rm G,D}$ is only weakly dependent on orbital radius, $M_{p,1}$ retains a similar scaling with $\dot{m}$. Thus, we see that the innermost planet’s mass varies from $M_{p,1}=1.18M_{\oplus}$ for $\dot{m}=10^{-10}\>M_{\odot}\>{\rm yr}^{-1}$, to $4.25M_{\oplus}$ for $\dot{m}=10^{-9}\>M_{\odot}\>{\rm yr}^{-1}$, to $15.3M_{\oplus}$ for $\dot{m}=10^{-8}\>M_{\odot}\>{\rm yr}^{-1}$ and that the required gap opening mass is significantly larger in the earlier, higher accretion rate phases when the DZIB is further out.
@@
$t_{\rm form,1}$ $=$ $\frac{M_{p,1}}{f_{\rm trap}\dot{m}_{p}}$
$=$ $1.28\times 10^{6}\frac{\gamma_{1.4}^{35/36}\kappa_{10}^{5/18}}{f_{\rm trap}f_{p}f_{s,0.01}}\alpha_{-4}^{2/9}m_{*,1}^{1/6}\frac{f_{r}^{5/9}}{\dot{m}_{-9}^{4/9}}\>{\rm yr}$
Thus, we see that even when the trapping fraction rises to unity the timescale to form the first planet can be $\sim 1\>$Myr, which is a characteristic timescale of disk accretion rate evolution. However, as discussed in Paper IV, there could be a phase of boosted pebble supply rate during early phases of disk evolution, triggered by the time when a significant population of macroscopic pebbles develops that is able to engage in efficient sweep-up growth of the small grain dust population. Typical boost factors in the pebble supply rate in the first few $\times 10^{5}\>$yr were found to be $f_{p,{\rm boost}}\sim 10$. In the case of significant enhancements in pebble supply rates, the timescales implied by equation (LABEL:eq:tform) should be regarded as upper limits.
@@
With an accretion rate that declines by an order of magnitude per Myr, both $r_{\rm DZIB}$ and the gap-opening mass $M_{p,1}$ evolve during the interval over which the planet accumulates most of its mass. Figure 5 indicates (solid circles for $f_{p,{\rm boost}}=1$; open circles for $f_{p,{\rm boost}}=10$) the locations along the evolutionary tracks when the time integrated trapped pebble mass equals the required planetary gap opening mass at the current DZIB location.
@@
The results discussed above may be influenced by the assumed evolution of the disk accretion rate, motivating a brief consideration of alternative $\dot{m}(t)$ histories. Our fiducial exponential decay model, in which $\dot{m}$ decreases by an order of magnitude per Myr, is motivated by Paper IV and is broadly consistent with observationally inferred disk lifetimes of $\sim 1$$10\ {\rm Myr}$. A slower decline in $\dot{m}$ would keep the DZIB at larger radii for longer and increase the associated gap-opening mass, thereby lengthening the formation time of the Vulcan planet and producing a more massive planetary system with generally wider orbital spacing. In more extreme cases, the system may remain in a high-accretion, low-trapping-efficiency state for an extended period, effectively stalling the onset of IOPF and yielding longer formation times and fewer planets. Conversely, a more rapid decay of $\dot{m}$ or early depletion of the inward pebble flux could inhibit the growth of the first planet before it reaches the shallow gap-opening mass to trigger DZIB retreat, thereby halting the subsequent IOPF process. However, it remains uncertain whether outer dust traps can form before the development of the DZIB. Rings and gaps are ubiquitous features in gaseous disks, yet their formation mechanisms—and consequently their characteristic timescales—are still not well constrained. Moreover, recent theoretical studies suggest that these substructures behave more like dust “traffic jams” rather than efficient dust traps (e.g., Hu et al., 2022). Such rapid $\dot{m}$ declines are also in tension with viscous evolution models, as they imply unphysically fast depletion of the gas disk mass. While wind-driven accretion can remove gas more efficiently, it does not substantially alter the midplane dust reservoir, and a detailed examination of pebble delivery in wind-driven disks is beyond the scope of this work.